Cost of common stock equity-CAPM J&M Corporation common stock has a beta, b, of 0.6. The risk-free rate is 4%, and the market return is 9% a. Determine the risk premium on J&M common stock. b. Determine the required return that J&M common stock should provide. c. Determine J&Ms cost of common stock equity using the CAPM.
Answer #1

a.Risk premium=Market rate-Risk free rate

=(9-4)=5%

b.Required return=Risk free rate+Beta*Risk premium

=4+0.6*5=7%

c.cost of common stock= risk-free rate +Beta*(MArket rate- risk-free rate )

=4+0.6*(9-4)=7%